# An Order of Magnitude Time Complexity Reduction for Gaussian Graphical Model Posterior Sampling Using a Reverse Telescoping Block Decomposition

### v1

@@ Acknowledgments
Nest, we consider the following equivalent representations of the partial likelihood:
@@
Standard calculations now yield that the posterior be $\boldsymbol{\beta}_{j}$ is:
@@
    Hubs. The $p$ rows and columns are partitioned into $K$ disjoint groups $\{\mathbf{G}_{k}\}_{k=1}^{K}$. For $i\neq j$, the off-diagonal entry is set to $0.25$ if $i$ and $j$ belong to the same group, and 0 otherwise. In our simulations, we set $K=p/10$, with each group containing the same number of nonzero elements.
@@
    Cliques negative. The $p$ rows and columns are partitioned into $K$ disjoint groups. We select 3 members within each group, $g_{k}\subset\mathbf{G}_{k}$. the off-diagonal entry is set to $0.75$ (corresponding to negative partial correlations) if $i\neq j$ and $i,j\in g_{k}$, and 0 otherwise. we set $K=p/10$.
@@
Table S.1: Estimation results for the RT and cyclical samplers under the tridiagonal structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
25% 10.93(0.27) 10.74(0.22) 3.58(0.08) 3.51(0.06) 0.60(0.17)
50% 12.11(0.32) 11.90(0.26) 3.74(0.13) 3.64(0.10) 0.52(0.15)
75% 13.50(0.37) 13.24(0.30) 4.42(0.24) 4.23(0.19) 0.62(0.14)
(127, 200) Mean 17.30(0.33) 16.84(0.27) 5.03(0.18) 4.80(0.12) 0.84(0.12)
Var 0.55(0.03) 0.50(0.02) 14.59(0.02) 14.61(0.01) 0.12(0.01)
25% 15.58(0.30) 15.19(0.24) 5.06(0.12) 4.96(0.09) 1.04(0.17)
50% 17.16(0.33) 16.71(0.27) 5.28(0.15) 5.11(0.10) 1.01(0.14)
75% 19.00(0.38) 18.48(0.31) 6.17(0.24) 5.83(0.17) 1.14(0.15)
(137, 300) Mean 21.08(0.33) - 6.13(0.16) - -
Var 0.64(0.03) - 17.90(0.02) - -
25% 19.06(0.29) - 6.17(0.12) - -
50% 20.93(0.33) - 6.46(0.14) - -
75% 23.10(0.37) - 7.49(0.22) - -
Table S.2: Estimation results for the RT and cyclical samplers under the hubs structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 15.02(0.51) 14.34(0.39) 5.88(0.52) 5.19(0.39) 1.09(0.20)
Var 0.61(0.03) 0.53(0.02) 10.10(0.02) 10.14(0.01) 0.14(0.02)
25% 13.42(0.46) 12.83(0.35) 5.01(0.40) 4.51(0.28) 1.23(0.21)
50% 14.89(0.50) 14.23(0.39) 5.78(0.49) 5.16(0.37) 1.08(0.21)
75% 16.62(0.56) 15.85(0.42) 7.22(0.58) 6.41(0.44) 1.23(0.24)
(127, 200) Mean 21.53(0.58) 20.31(0.44) 8.51(0.57) 7.30(0.43) 1.77(0.24)
Var 0.78(0.04) 0.66(0.02) 14.35(0.02) 14.41(0.01) 0.22(0.03)
25% 19.38(0.53) 18.31(0.41) 7.26(0.46) 6.34(0.32) 1.94(0.25)
50% 21.36(0.57) 20.19(0.44) 8.40(0.54) 7.30(0.40) 1.78(0.24)
75% 23.67(0.63) 22.31(0.48) 10.33(0.64) 8.94(0.48) 2.06(0.27)
(137, 300) Mean 26.55(0.57) - 10.53(0.60) - -
Var 0.91(0.03) - 17.61(0.02) - -
25% 23.98(0.53) - 8.97(0.47) - -
50% 26.36(0.57) - 10.39(0.57) - -
75% 29.10(0.61) - 12.71(0.66) - -
Table S.3: Estimation results for the RT and cyclical samplers under the cliques negative structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 25.27(1.14) 23.82(1.08) 15.77(1.22) 14.25(1.18) 1.69(0.30)
Var 1.46(0.15) 1.31(0.13) 10.53(0.08) 10.62(0.07) 0.23(0.08)
25% 22.64(1.02) 21.32(0.97) 13.29(1.10) 11.92(1.06) 1.59(0.29)
50% 25.08(1.13) 23.65(1.08) 15.58(1.22) 14.09(1.17) 1.65(0.31)
75% 27.70(1.24) 26.15(1.19) 18.10(1.33) 16.48(1.28) 1.84(0.36)
(127, 200) Mean 35.66(1.13) 33.71(1.08) 22.17(1.21) 20.17(1.17) 2.25(0.23)
Var 1.77(0.12) 1.57(0.11) 15.09(0.07) 15.21(0.06) 0.30(0.06)
25% 32.22(1.03) 30.45(0.98) 18.93(1.10) 17.12(1.06) 2.16(0.25)
50% 35.42(1.12) 33.50(1.08) 21.94(1.20) 19.97(1.16) 2.20(0.23)
75% 38.86(1.22) 36.77(1.17) 25.25(1.31) 23.09(1.26) 2.46(0.27)
(137, 300) Mean 43.25(0.98) - 26.68(1.05) - -
Var 1.96(0.10) - 18.62(0.06) - -
25% 39.24(0.90) - 22.92(0.96) - -
50% 42.98(0.98) - 26.42(1.05) - -
75% 46.99(1.06) - 30.27(1.13) - -
Figure S.1: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the tridiagonal structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.2: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the hubs structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.3: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the cliques negative structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
@@
Table S.4: Estimation results for the RT and cyclical samplers under the tridiagonal structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 11.69(0.36) 12.01(0.25) 3.39(0.14) 3.44(0.12) 0.61(0.19)
Var 0.40(0.03) 0.39(0.02) 10.29(0.03) 10.29(0.01) 0.09(0.02)
25% 10.43(0.30) 10.75(0.22) 3.51(0.06) 3.51(0.06) 0.83(0.21)
50% 11.59(0.35) 11.91(0.25) 3.64(0.11) 3.64(0.10) 0.78(0.24)
75% 12.92(0.42) 13.25(0.29) 4.18(0.21) 4.23(0.19) 0.93(0.33)
(127, 200) Mean 16.81(0.47) 16.95(0.26) 4.85(0.17) 4.84(0.12) 0.71(0.12)
Var 0.53(0.03) 0.51(0.02) 14.60(0.02) 14.61(0.01) 0.10(0.01)
25% 15.13(0.41) 15.29(0.23) 4.99(0.09) 4.98(0.09) 0.95(0.16)
50% 16.68(0.46) 16.83(0.26) 5.16(0.14) 5.13(0.11) 0.91(0.20)
75% 18.47(0.53) 18.61(0.29) 5.92(0.27) 5.89(0.16) 1.06(0.22)
(137, 300) Mean 21.04(0.33) - 6.11(0.16) - -
Var 0.64(0.03) - 17.90(0.02) - -
25% 19.00(0.29) - 6.16(0.11) - -
50% 20.88(0.33) - 6.44(0.14) - -
75% 23.06(0.38) - 7.47(0.22) - -
Table S.5: Estimation results for the RT and cyclical samplers under the hubs structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 14.73(0.52) 14.32(0.38) 5.62(0.51) 5.16(0.38) 0.86(0.17)
Var 0.62(0.04) 0.53(0.02) 10.09(0.02) 10.14(0.01) 0.15(0.02)
25% 13.10(0.46) 12.81(0.35) 4.78(0.38) 4.50(0.28) 1.01(0.17)
50% 14.60(0.51) 14.21(0.39) 5.57(0.48) 5.14(0.37) 0.87(0.19)
75% 16.34(0.57) 15.83(0.42) 6.98(0.58) 6.39(0.44) 1.00(0.21)
(127, 200) Mean 21.55(0.58) 20.34(0.41) 8.54(0.58) 7.33(0.41) 1.76(0.27)
Var 0.83(0.04) 0.66(0.02) 14.31(0.03) 14.41(0.01) 0.27(0.03)
25% 19.33(0.53) 18.34(0.38) 7.21(0.46) 6.37(0.31) 1.87(0.26)
50% 21.37(0.58) 20.21(0.41) 8.41(0.54) 7.32(0.38) 1.75(0.26)
75% 23.75(0.64) 22.34(0.44) 10.42(0.64) 8.97(0.46) 2.09(0.32)
(137, 300) Mean 26.26(0.49) - 10.27(0.52) - -
Var 0.97(0.03) - 17.57(0.02) - -
25% 23.61(0.44) - 8.65(0.40) - -
50% 26.06(0.48) - 10.14(0.50) - -
75% 28.87(0.53) - 12.51(0.57) - -
Table S.6: Estimation results for the RT and cyclical samplers under the cliques negative structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (“–”) indicates the cyclical sampler exceeded the 44-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 24.98(1.14) 23.83(1.09) 15.51(1.22) 14.27(1.18) 1.42(0.25)
Var 1.45(0.15) 1.30(0.13) 10.54(0.08) 10.63(0.07) 0.21(0.09)
25% 22.37(1.02) 21.34(0.98) 13.05(1.10) 11.93(1.06) 1.33(0.23)
50% 24.79(1.13) 23.67(1.08) 15.33(1.21) 14.10(1.17) 1.39(0.26)
75% 27.40(1.24) 26.16(1.19) 17.83(1.33) 16.49(1.28) 1.55(0.31)
(127, 200) Mean 35.66(1.09) 33.78(1.06) 22.17(1.17) 20.23(1.14) 2.21(0.31)
Var 1.80(0.12) 1.57(0.10) 15.07(0.07) 15.21(0.06) 0.34(0.07)
25% 32.19(0.99) 30.52(0.96) 18.91(1.06) 17.18(1.04) 2.07(0.30)
50% 35.41(1.08) 33.57(1.05) 21.93(1.16) 20.03(1.14) 2.14(0.30)
75% 38.88(1.18) 36.83(1.14) 25.27(1.26) 23.15(1.24) 2.44(0.35)
(137, 300) Mean 42.88(1.02) - 26.35(1.09) - -
Var 1.97(0.11) - 18.61(0.06) - -
25% 38.86(0.92) - 22.57(0.99) - -
50% 42.61(1.01) - 26.09(1.08) - -
75% 46.63(1.10) - 29.94(1.18) - -
Figure S.4: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the tridiagonal structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.5: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the hubs structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.6: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the cliques negative structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,150,200,250,300\}$, with $n\approx 24\log(p)$. At $p=300$, the cyclical sampler exceeded the 4-hour wall time limit, so its raw runtime is capped at 4 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.

### v2

@@ Acknowledgments
Next, we consider the following equivalent representations of the partial likelihood:
@@
Standard calculations now yield that the posterior of $\boldsymbol{\beta}_{j}$ is:
@@
    Cliques negative. The $p$ rows and columns are partitioned into $K$ disjoint groups. We select 3 members within each group, $g_{k}\subset\mathcal{G}_{k}$. The off-diagonal entry is set to $0.35$ (corresponding to negative partial correlations) if $i\neq j$ and $i,j\in g_{k}$, and 0 otherwise. We set $K=p/50$.
@@
    Cliques positive. The $p$ rows and columns are partitioned into $K$ disjoint groups. We select 3 members within each group, $g_{k}\subset\mathcal{G}_{k}$. The off-diagonal entry is set to 0.35-0.35 (corresponding to positive partial correlations) if $i\neq j$ and $i,j\in g_{k}$, and 0 otherwise. We set $K=p/50$.
@@
Table S.1: Estimation results for the RT and Cyclical samplers under the tridiagonal structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
25% 10.95(0.28) 10.77(0.23) 3.58(0.08) 3.52(0.06) 0.67(0.15)
50% 12.13(0.31) 11.92(0.26) 3.75(0.13) 3.65(0.11) 0.59(0.14)
75% 13.52(0.37) 13.26(0.31) 4.43(0.25) 4.24(0.19) 0.70(0.15)
(137, 300) Mean 21.07(0.35) 20.62(0.29) 6.13(0.19) 5.89(0.12) 1.03(0.15)
Var 0.64(0.02) 0.59(0.02) 17.90(0.02) 17.93(0.01) 0.15(0.01)
25% 19.05(0.32) 18.67(0.26) 6.17(0.14) 6.05(0.09) 1.36(0.17)
50% 20.91(0.35) 20.48(0.29) 6.45(0.17) 6.25(0.10) 1.28(0.17)
75% 23.07(0.39) 22.56(0.33) 7.48(0.25) 7.13(0.18) 1.47(0.18)
(160, 800) Mean 34.46(0.46) - 10.07(0.20) - -
Var 0.96(0.03) - 29.32(0.02) - -
25% 31.36(0.41) - 10.06(0.12) - -
50% 34.24(0.45) - 10.57(0.17) - -
75% 37.56(0.52) - 12.20(0.29) - -
Table S.2: Estimation results for the RT and Cyclical samplers under the cliques negative structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 10.42(0.11) 10.42(0.09) 1.23(0.10) 1.24(0.09) 0.33(0.26)
Var 0.22(0.03) 0.22(0.02) 9.87(0.01) 9.88(0.01) 0.06(0.04)
25% 9.44(0.09) 9.44(0.08) 1.36(0.09) 1.37(0.09) 0.29(0.27)
50% 10.35(0.11) 10.35(0.09) 1.25(0.09) 1.26(0.07) 0.31(0.33)
75% 11.34(0.13) 11.33(0.11) 1.78(0.12) 1.79(0.11) 0.48(0.40)
(137, 300) Mean 18.13(0.14) 18.02(0.12) 2.15(0.16) 2.14(0.10) 0.87(0.36)
Var 0.33(0.02) 0.32(0.02) 17.16(0.01) 17.17(0.01) 0.11(0.03)
25% 16.59(0.12) 16.50(0.10) 2.31(0.12) 2.33(0.08) 0.91(0.36)
50% 18.03(0.14) 17.93(0.12) 2.18(0.15) 2.18(0.10) 1.02(0.38)
75% 19.58(0.16) 19.45(0.14) 3.03(0.18) 2.97(0.14) 1.16(0.45)
(160, 800) Mean 29.96(0.18) - 3.60(0.17) - -
Var 0.53(0.02) - 28.07(0.01) - -
25% 27.58(0.16) - 3.62(0.14) - -
50% 29.81(0.18) - 3.59(0.14) - -
75% 32.22(0.21) - 5.10(0.21) - -
Table S.3: Estimation results for the RT and Cyclical samplers under the cliques positive structure for the GHS prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 11.09(0.13) 11.15(0.16) 2.15(0.31) 2.23(0.34) 0.74(0.39)
Var 0.33(0.06) 0.32(0.05) 9.88(0.01) 9.88(0.01) 0.14(0.06)
25% 10.19(0.14) 10.24(0.16) 2.17(0.30) 2.24(0.33) 0.94(0.49)
50% 11.03(0.13) 11.08(0.16) 2.14(0.31) 2.22(0.33) 0.82(0.52)
75% 11.97(0.13) 12.02(0.16) 2.68(0.25) 2.75(0.28) 0.87(0.42)
(137, 300) Mean 19.20(0.17) 19.22(0.19) 3.74(0.36) 3.85(0.39) 1.17(0.32)
Var 0.47(0.05) 0.46(0.05) 17.17(0.01) 17.18(0.01) 0.21(0.05)
25% 17.79(0.17) 17.82(0.19) 3.72(0.35) 3.84(0.39) 1.37(0.39)
50% 19.10(0.17) 19.12(0.18) 3.74(0.37) 3.84(0.39) 1.37(0.45)
75% 20.57(0.17) 20.58(0.19) 4.54(0.32) 4.60(0.34) 1.45(0.32)
(160, 800) Mean 31.55(0.16) - 6.09(0.36) - -
Var 0.68(0.04) - 28.09(0.01) - -
25% 29.37(0.16) - 5.92(0.37) - -
50% 31.39(0.16) - 6.04(0.36) - -
75% 33.66(0.17) - 7.39(0.29) - -
Figure S.1: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the tridiagonal structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.2: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the cliques negative structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.3: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHS prior under the cliques positive structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
@@
Table S.4: Estimation results for the RT and Cyclical samplers under the tridiagonal structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on up to $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 11.85(0.60) 12.00(0.27) 3.49(0.16) 3.44(0.13) 0.76(0.65)
Var 0.40(0.06) 0.39(0.02) 10.29(0.04) 10.29(0.01) 0.11(0.06)
25% 10.59(0.50) 10.75(0.24) 3.55(0.08) 3.52(0.06) 0.86(0.49)
50% 11.74(0.58) 11.90(0.27) 3.68(0.13) 3.64(0.10) 0.86(0.65)
75% 13.08(0.71) 13.24(0.31) 4.27(0.28) 4.23(0.19) 1.03(0.87)
(137, 300) Mean 20.46(1.13) 20.61(0.30) 6.08(0.30) 5.89(0.13) 1.63(0.93)
Var 0.59(0.10) 0.58(0.02) 17.93(0.07) 17.93(0.01) 0.21(0.08)
25% 18.51(0.97) 18.66(0.27) 6.16(0.16) 6.05(0.09) 1.85(0.63)
50% 20.31(1.10) 20.47(0.30) 6.39(0.24) 6.26(0.11) 1.93(0.89)
75% 22.38(1.30) 22.55(0.34) 7.29(0.51) 7.13(0.18) 2.31(1.25)
(160, 800) Mean 33.49(1.63) - 9.91(0.51) - -
Var 0.90(0.13) - 29.36(0.08) - -
25% 30.50(1.42) - 10.00(0.29) - -
50% 33.29(1.60) - 10.41(0.45) - -
75% 36.46(1.86) - 11.81(0.83) - -
Table S.5: Estimation results for the RT and Cyclical samplers under the cliques negative structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on up to $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 10.37(0.11) 10.39(0.08) 1.24(0.08) 1.24(0.08) 0.40(0.35)
Var 0.21(0.03) 0.21(0.01) 9.88(0.01) 9.88(0.00) 0.06(0.04)
25% 9.40(0.09) 9.42(0.07) 1.37(0.07) 1.38(0.08) 0.36(0.34)
50% 10.31(0.10) 10.33(0.07) 1.25(0.07) 1.26(0.09) 0.35(0.39)
75% 11.28(0.13) 11.30(0.09) 1.75(0.10) 1.77(0.09) 0.51(0.48)
(137, 300) Mean 18.07(0.32) 18.02(0.13) 2.20(0.20) 2.15(0.12) 0.92(0.35)
Var 0.32(0.04) 0.32(0.02) 17.17(0.02) 17.17(0.01) 0.13(0.04)
25% 16.55(0.28) 16.50(0.12) 2.34(0.11) 2.34(0.10) 0.96(0.36)
50% 17.98(0.30) 17.92(0.13) 2.20(0.17) 2.18(0.11) 0.96(0.39)
75% 19.51(0.36) 19.45(0.15) 3.03(0.30) 2.96(0.15) 1.18(0.43)
(160, 800) Mean 29.77(0.64) - 3.60(0.38) - -
Var 0.50(0.08) - 28.08(0.03) - -
25% 27.42(0.56) - 3.68(0.18) - -
50% 29.62(0.61) - 3.58(0.29) - -
75% 32.00(0.71) - 4.98(0.58) - -
Table S.6: Estimation results for the RT and Cyclical samplers under the cliques positive structure for the GHSL prior. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are based on up to $50$ replicates and Monte Carlo standard errors are in parentheses. A dash (–) indicates the cyclical sampler exceeded the $12$-hour runtime limit and no results were obtained.
@@
(110, 100) Mean 11.02(0.17) 11.13(0.15) 2.16(0.31) 2.21(0.34) 0.93(0.50)
Var 0.28(0.06) 0.33(0.06) 9.88(0.01) 9.88(0.01) 0.16(0.07)
25% 10.11(0.18) 10.23(0.15) 2.19(0.31) 2.24(0.34) 1.12(0.59)
50% 10.95(0.17) 11.06(0.15) 2.14(0.30) 2.21(0.34) 0.95(0.58)
75% 11.89(0.17) 12.00(0.15) 2.62(0.27) 2.74(0.27) 1.00(0.48)
(137, 300) Mean 19.15(0.31) 19.22(0.19) 3.81(0.37) 3.84(0.38) 1.43(0.46)
Var 0.42(0.07) 0.46(0.05) 17.17(0.01) 17.18(0.01) 0.24(0.06)
25% 17.74(0.29) 17.82(0.19) 3.79(0.36) 3.84(0.38) 1.73(0.53)
50% 19.05(0.30) 19.12(0.19) 3.78(0.36) 3.84(0.37) 1.53(0.52)
75% 20.52(0.34) 20.58(0.19) 4.52(0.37) 4.59(0.34) 1.61(0.41)
(160, 800) Mean 31.38(0.62) - 6.08(0.44) - -
Var 0.64(0.09) - 28.10(0.02) - -
25% 29.21(0.56) - 5.95(0.35) - -
50% 31.23(0.60) - 6.02(0.39) - -
75% 33.48(0.68) - 7.26(0.57) - -
Figure S.4: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the tridiagonal structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.5: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the cliques negative structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.
Figure S.6: Average raw runtimes in seconds (left panel) and runtimes relative to $p=100$ (right panel), for the GHSL prior under the cliques positive structure. Each MCMC run uses 10,00010{,}000 iterations with the first 5,0005{,}000 discarded as burn-in. Results are averaged over 50 replications for each $p\in\{100,200,300,400,600,800\}$, with $n=24\log p$. At $p>400$, the cyclical sampler exceeded the 12-hour wall time limit, so its raw runtime is capped at 12 hours. The RT sampler is implemented in Rcpp and the Cyclical sampler in R.